OpenStudy (monkey*):

What is equivalent to 3^4×9^3 3^4*(3^2)^3 3^4*3^6 None

OpenStudy (maddy1251):

Do you know what 3^4 = ? What does9 ^ 3 = ?

OpenStudy (maddy1251):

3 x 3 x 3 x 3 = ? (3 x 3 ) x (3 x 3) = ? (9) x (9) = ? This makes it easier

OpenStudy (maddy1251):

And then you can do (9 x 9 x 9) = ? (9x9) x (9) = ? (81) x 9 =?

OpenStudy (maddy1251):

3^4 = 81 9^3 = 729, fyi.

OpenStudy (maddy1251):

Hope that part makes sense. Just break down the numbers and group them to make it the easiest it can possibly be!

OpenStudy (maddy1251):

3^4*(3^2)^3 So, we know (3 x 3 x 3 x 3) = 81. (3^2)= 3 x 3 = 9 NOW we do 9^3 = 729. Looks familiar.. :) Lets move onto the next option

OpenStudy (monkey*):

You're making it easier to understand, thank you!

OpenStudy (maddy1251):

3^4*3^6 3^4 = (3 x 3 x 3 x 3) = 81 (This is a very popular number!) (3 x 3 x 3 x 3 x 3 x 3) = v (3 x 3 ) x (3 x 3) x (3x3) = 9 x 9 x 9 = 729

OpenStudy (maddy1251):

Not sure if you can pick two answers, however each seem to equal out to be the same. I'm gonna make sure I didn't mess up anywhere!

OpenStudy (maddy1251):

They probably want choice A, and I'll explain why!

OpenStudy (monkey*):

You can pick two!

OpenStudy (maddy1251):

Oh well then I would say both, from what I worked out then :) Do you agree?

OpenStudy (monkey*):

Yes, I agree with you!

OpenStudy (maddy1251):

Choice A technically yields : 3^4*(3^2)^3 The 3^4 is just the same as the original (3^2) = 9 Then goes 9^3, just the same as before! However both equal the same end result of 59049

OpenStudy (maddy1251):

So I would personally pick both. :-)

OpenStudy (monkey*):

Do you mind helping with a few more? I would be totally grateful!

OpenStudy (maddy1251):

Sure thing! Anything to help out

OpenStudy (monkey*):

(5^3*5^2)^4 25^20 (25^5)^4 None

OpenStudy (maddy1251):

Do you know how to work it out? Or where to start?

OpenStudy (monkey*):

5*5*5 5*5 ?

OpenStudy (maddy1251):

Yep! Within the parentheses is ((5x5x5)x(5x5)) = (5x5x5x5x5)^4 Keep going! Lets forget about that pesky 4. What is 5^5?

OpenStudy (monkey*):

5^3=125 5^2=25

OpenStudy (maddy1251):

Multiply them together :)

OpenStudy (monkey*):

3125?

OpenStudy (maddy1251):

Yes! We now know what (5^3*(5^2)) equals! It is now 3125 ^ 4. What does that equal?

OpenStudy (monkey*):

12500

OpenStudy (maddy1251):

Not quite! It is actually going to equal 95367431640625. Big number, I know! We are not taking 3125 x 4 rather we are doing... 3125 x 3125 x 3125 x 3125

OpenStudy (monkey*):

I did that, but wasn't sure if that was right!

OpenStudy (maddy1251):

Now, 25^20 = ? (25^5)^4 = ? Well, we can simply punch in 25^20 into the calculator. That would be typing 25 too many times. 25^20 = 9094947017729282379150390625 Then lets do (25^5)^4 = ? Well.. (25^5) = 9765625 = (9765625)^4 = 9094947017729282379150390625 Are these related to 3125^4?

OpenStudy (monkey*):

Doesn't seem like they are the same.

OpenStudy (maddy1251):

You are correct! They are way too big compared to 95367431640625 :-)

OpenStudy (monkey*):

So the answer would be none! Right?(:

OpenStudy (maddy1251):

that's what I would go with!

OpenStudy (monkey*):

It was right!

OpenStudy (maddy1251):

Awesome! Do you understand how we got the answer?

OpenStudy (monkey*):

Yes I do. You've helped so much, thank you so much!

OpenStudy (maddy1251):

You are very much welcome :)

OpenStudy (mathstudent55):

I think the idea with these problems is to make you practice the rules of operations with exponents. There is no need to multiply out the powers and get enormous numbers. You need to apply the rules of exponents and see if the expressions are equal.

OpenStudy (mathstudent55):

The first problem: 3^4×9^3 A. 3^4*(3^2)^3 B. 3^4*3^6 C. None Since 9 = 3^2, then 9^3 = (3^2)^2 That means that 3^4×9^3 = 3^4×(3^2)^3, and choice A is true Since \((a^m)^n = a^{mn}\), then (3^2)^3 = 3^(2*3) = 3^6, so 3^4×9^3 = 3^4×(3^2)^3 = 3^4×3^6, and choice B is also true.

OpenStudy (mathstudent55):

The second problem: (5^3*5^2)^4 A.25^20 B. (25^5)^4 C. None Since 5^3 * 5^2 = 5^(3 + 2) = 5^5, then (5^3*5^2)^4 = (5^5)^4 = 5^20, and choice A is true. Let's work on choice B. 25 = 5^2, so (25^5)^4 = 25^20 = (5^2)^20 = 5^40 Earlier we saw the expression is equal to 5^20, so choice B is not true.